log[3](0.1)+2log[3](x)=log[3](2)+log[3](5)solve the equation if necessary round to four decimal places
Answer by stanbon(75887)

log[3](0.1)+2log[3](x)=log[3](2)+log[3](5)solve the equation if necessary round to four decimal places
--------
= log3[0.1*x^2] = log3[2*5]
--------
0.1*x^2 = 10
---
x^2 = 100
---
x = 10
===============
Cheers,
Stan H.
==================